what is FIR filter and IIR filter? (include frequency response and z transform)
Answer #1

FIR filter:

In this filter output is weighted sum of current and previous M inputs.

Filter has order of M

Impulse response is of length M+1

Difference equation is

y(n) = \sum_{k=0}^{M}b_{k}x(n-k)

if input is complex Sinusoid

y(n) = \sum_{k=0}^{M}b_{k} Ae^{j\phi }e^{j\omega (n-k)}

y(n) -Aeloej b.eji eJeJa k e

jpjw(n)

Frequency response of FIR filter is

H(e^{j\omega})= \sum_{k=0}^{M}b_{k} e^{-j\omega k}

Similarly in z-transform also

に0

IIR filter:

Output is weighted sum of current and previous M inputs and previous N outputs

.V k=1 k=0

Frequency response of IIR filter can be found similar to FIR filter

.V Y(z) = Σ ak Y(z):-k + Σ bks@:-k k=1 に0

\frac{Y(z)}{X(z)} =\frac{ \sum_{k=0}^{M}b_{k} z^{-k}}{1 - \sum_{k=1}^{N}a_{k} z^{-k}}

1- A

For frequency response we replace

H(e^{j\omega}) = H(z)|_{z=e^{j\omega}}

Hiei)- bre-juk d 1-